OpenStudy (anonymous):

One quick algebra question pleaseee!!!! medal rewarded!!! Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (solomonzelman):

\[\sqrt{2} \times (5 + \sqrt{8})\] you can exclude A and B right away, because, \[5\sqrt{2}\]is irrational. Solve it, \[5\sqrt{2}+4\] so which option is the answer?

OpenStudy (anonymous):

So it's D, correct?

OpenStudy (solomonzelman):

YUP!

OpenStudy (anonymous):

thank you so much! (:

OpenStudy (solomonzelman):

You Welcome!